# Persistent Charge and Spin Currents in a Ferromagnetic Hatano-Nelson Ring

### v1

@@ I Introduction
Despite these recent advances, further efforts are needed to explore persistent currents in non-Hermitian systems. In particular, to the best of our knowledge, persistent spin currents in non-Hermitian systems have not been investigated, and no theoretical framework has yet been established to describe their behavior. Understanding spin-dependent currents in this context is therefore of both fundamental interest and potential practical relevance. Therefore, in the present work, we focus on electronic rings with the same type of non-Hermiticity previously employed for persistent charge currents [39], providing the first systematic study of persistent spin current behavior.
Specifically, we study a dimerized HN ring with anti-Hermitian intradimer hopping [39], incorporating a uniform ferromagnetic ordering to explore spin-resolved persistent currents and their interplay with non-Hermiticity. The system is modeled within the nearest-neighbor tight-binding framework. Persistent charge and spin currents in mesoscopic rings can be computed through different approaches, including the derivative method [42, 2], the current operator formalism [26, 30], and Green’s function techniques [29]. While the derivative method depends only on the eigenvalues of the Hamiltonian, both the current operator and Green’s function approaches require explicit knowledge of the eigenstates. In this work, we adopt the current operator formalism.
@@ II.1 Schematic and model Hamiltonian
II.2 Dispersion relations
By writing the Hamiltonian in $k$-space in the basis
@@ II.2 Dispersion relations
where the superscript indicates bonding (-) or antibonding (++) and the subscript indicates spin orientation along $\mathbf{m}$. The full derivation, starting from the Bloch Hamiltonian and proceeding up to the energy spectrum, is provided in Appendix A.
@@
In the upper panel (Figs. 3(a) and (b)), corresponding to $\lvert t_{1}\rvert=0.75$, the system lies in the topological phase, analogous to the non-Hermitian Hatano-Nelson model [36, 39]. The real
part of the spectrum (Fig. 3(a)) exhibits a clear bulk energy gap, characteristic of a topologically nontrivial band structure. The imaginary part (Fig. 3(b)), however, is gapless, the bands are smoothly connected across the Brillouin zone without any separation in their imaginary components.
@@
II.3 Persistent charge current
The charge current operator is defined as
$\hat{J}=\frac{e\dot{X}}{2Na},\quad\dot{X}=\frac{2\pi}{ih}[X,H],$ (3)
where $e$ is the electronic charge, $a$ is the lattice spacing, and $2N$ is the total number of sites. $X$ and $\dot{X}$ are the position and velocity operators, respectively. In the non-Hermitian case, the current for the $n$th eigenstate is evaluated in the biorthogonal basis (a discussion of biorthogonal basis is given in Appendix B) through the operation as
where $\lvert\psi_{n}^{R}\rangle$ and $\lvert\psi_{n}^{L}\rangle$ denote the right and left eigenvectors, respectively. The explicit derivation of $\hat{J}$ and its evaluation in terms of the expansion coefficients of $\lvert\psi_{n}^{R/L}\rangle$ is provided in Appendix C.
II.4 Persistent spin current
The spin current operator is defined as
The explicit operator forms of $I_{x}$, $I_{y}$, and $I_{z}$, together with their evaluation in terms of the expansion coefficients of $\lvert\psi_{n}^{R/L}\rangle$, are provided in Appendix D.
III Results and Discussion
[…]
To gain some physical understanding of the results, we begin by analyzing a simplified setup. Specifically, we consider the case where all magnetic moments are aligned along the $z$-quantization axis, i.e., $\theta=0$ and $\phi=0$. In this limit, the Hamiltonian (Eq. 1) decouples into independent up- and down-spin sectors, $\left(H_{\uparrow}+H_{\downarrow}\right)$. We set the SDS parameter to $\mathpzc{h}=0.5$ and fix the number of unit cells in the Hatano-Nelson ring at $N=8$. The resulting complex energy spectra for the up- and down-spin Hamiltonians are shown as a function of $\Phi$ in Fig. 5, for three representative values of the intradimer hopping amplitude. Since the hopping amplitudes follow an SSH-type dimerization, we examine three characteristic regimes: the topological phase $\lvert t_{1}\rvert<t_{2}$, the critical point $\lvert t_{1}\rvert=t_{2}$, and the trivial phase $\lvert t_{1}\rvert>t_{2}$.
@@ III Results and Discussion
Figure 6: (Color online.) Real and imaginary persistent charge current as a function of the normalized magnetic flux $\Phi/\pi$. The first, second, and third columns correspond to $t_{1}=0.75$, $1$, and $1.25$, respectively. The number of unit cells is $N=20$, the interdimer hopping amplitude is $t_{2}=1$, the spin-dependent splitting is $\mathpzc{h}=0.25$, and the chemical potential is $\mu=0$.
@@
Turning to the magnitudes, we observe the following. In the topological phase (Fig. 9(a)), all components are small, within $\pm 0.06\,$eV, reflecting the weak current response in this regime. At the critical point (Fig. 9(b)), the amplitudes of all three components grow significantly, with $I_{z}$ spanning nearly $\pm 0.5\,$eV. In the trivial phase (Fig. 9(c)), the overall trend resembles the earlier cases but with reduced amplitudes (about $\pm 0.15\,$eV) compared to the critical point. Importantly, the systematic behavior observed for the real components of the persistent charge current does not extend to the persistent spin currents. Specifically, while the charge currents are typically weak in the topological regime, moderate at the critical point, and largest in the trivial phase, the persistent spin currents do not follow this universal trend. Instead, their magnitudes are strongly dependent on the orientation of the magnetic moments, the chemical potential, and the SDS factor $\mathpzc{h}$.
@@
In terms of magnitudes, the topological phase (Fig. 10(a)) again shows weak response, in close analogy with the $\theta$-variation (Fig. 9(a)). Here $I_{z}$ is pinned near 0.030.03\,eV, while $I_{x}$ and $I_{y}$ oscillate within $\pm 0.05\,$eV. At the critical point (Fig. 10(b)), the magnitudes are strongly enhanced. $I_{z}$ remains fixed at about 0.2-0.2\,eV, and $I_{x}$ and $I_{y}$ extend almost up to $\pm 0.4\,$eV. In the trivial phase (Fig. 10(c)), the magnitudes are reduced once again, paralleling the $\theta$-variation (Fig. 9(c)). Here $I_{z}$ stays constant at about 0.060.06\,eV, while $I_{x}$ and $I_{y}$ oscillate within $\pm 0.12\,$eV. Taken together with the $\phi$-dependence and the $\theta$-variation, these results highlight that the spin persistent currents do not follow the universal pattern of the charge persistent current. Instead, they are strongly modulated by the orientation of the magnetic moments and respond anisotropically in spin space.
@@
For completeness, we also study the finite-size effect on the persistent spin currents by varying the number of unit cells from 10 to 60, while keeping the chemical potential fixed at $\mu=-0.5$. All other parameters are listed in the caption of Fig. 13, where we plot the three components of the persistent spin current as a function of the number of sites ($2N$, NN being the number of unit cells). We consider three different regimes defined by the intradimer hopping amplitude $t_{1}$, with the $x$, $y$, and $z$ components of the persistent spin current represented in red, blue, and green, respectively. Overall, the magnitudes of the components across all phases are qualitatively similar to the $\theta$- and $\varphi$-dependence discussed earlier. A systematic decrease of the current is observed with increasing system size for all components. However, the decay is fastest in the topological phase (Fig. 13(a)), where the current becomes negligibly small beyond $2N=40$. At the critical point (Fig. 13(b)), although the current still decreases with ring size, the rate of decay is much slower, and all components remain finite even at $2N=120$. In the trivial phase (Fig. 13(c)), the decay is moderate. This finite-size trend is fully consistent with our previous work [39], where the persistent charge current exhibits a similar behavior with system dimension.
@@ IV Summary
$\bullet$ The imaginary persistent charge current dominates in the topological phase, both components are enhanced at the critical point, and the real part dominates in the trivial phase, while the real and imaginary currents are no longer identical in the presence of the ferromagnetic exchange field.
@@
Appendix A Derivation of Bloch Hamiltonian and dispersion relations
[…]
We can write the Hamiltonian in block form as
$H(k)=\begin{pmatrix}A&B\mathbb{I}_{2}\\ B^{\prime}\mathbb{I}_{2}&A\end{pmatrix},\qquad D=\begin{pmatrix}0&B\\ B^{\prime}&0\end{pmatrix}$, (9)
where $\mathbb{I}_{2}$ is the $2\times 2$ identity matrix and
$B$ $=\lvert t_{1}\rvert e^{i\phi}+t_{2}e^{-ik},$ (10)
$B^{\prime}$ $=-\lvert t_{1}\rvert e^{-i\phi}+t_{2}e^{ik},$ (11)
[…]
Here $A$ acts in spin space and $D$ acts in sublattice space, so they commute. The Hamiltonian can therefore be expressed as
$H(k)=\mathbb{I}_{2}\otimes A+D\otimes\mathbb{I}_{2}.$ (13)
The eigenvalues of $A$ are
$\lambda_{A}=\epsilon\pm\mathpzc{h},$ (14)
and the eigenvalues of $D$ are
$\lambda_{D}=\pm\sqrt{BB^{\prime}},$ (15)
with
Explicitly,
Compactly, the eigenvalues can be written as
Appendix B Biorthogonalization of Eigenstates

### v2

@@ I Introduction
Although some progress has been made in understanding flux-sensitive responses in non-Hermitian lattice systems, the role of spin degrees of freedom in non-Hermitian persistent transport remains largely unexplored. While Hatano and Nelson did not explicitly formulate persistent currents in the mesoscopic sense, their demonstration of strong spectral sensitivity to boundary conditions and synthetic gauge phases in non-reciprocal lattices laid the foundation for later interpretations of current-carrying responses in non-Hermitian rings [33, 34].
In Hermitian systems, persistent spin currents are often closely connected to their charge counterparts and can frequently be understood within a unified framework. In contrast, non-Hermitian Hamiltonians fundamentally alter the structure of eigenstates and expectation values, necessitating a biorthogonal description in which spin-resolved observables acquire a distinct character. The interplay between non-reciprocal hopping and magnetic exchange therefore raises basic questions regarding the existence, structure, and physical interpretation of persistent spin currents in non-Hermitian systems.
Motivated by these open issues, we investigate persistent spin currents in a non-Hermitian dimerized ring with ferromagnetic exchange, employing the same class of asymmetric hopping previously used to study persistent charge currents [39]. The dimerized Hatano–Nelson ring provides a minimal and controlled platform in which non-reciprocity, topology, and magnetism can be tuned independently, allowing us to isolate the impact of spin degrees of freedom on persistent transport without introducing additional model-specific complexities. By systematically analyzing spin-resolved currents across different parameter regimes, we elucidate how non-Hermiticity and magnetism jointly shape persistent spin transport, thereby extending the understanding of equilibrium-like responses in non-Hermitian quantum systems beyond the spinless case.
Specifically, we study a dimerized HN ring with anti-Hermitian intradimer hopping [39], incorporating a uniform ferromagnetic ordering to explore spin-resolved persistent currents and their interplay with non-Hermiticity. The ferromagnetic configuration is chosen as the simplest magnetic background that breaks spin degeneracy while preserving translational symmetry, thereby allowing us to isolate the effects of non-Hermiticity on spin-dependent transport without additional complications arising from magnetic frustration or spatially varying textures. The system is modeled within the nearest-neighbor tight-binding framework. Persistent charge and spin currents in mesoscopic rings can be computed through different approaches, including the derivative method [42, 2], the current operator formalism [26, 30], and Green’s function techniques [29]. While the derivative method depends only on the eigenvalues of the Hamiltonian, both the current operator and Green’s function approaches require explicit knowledge of the eigenstates. In this work, we adopt the current operator formalism.
@@
Before proceeding, we briefly clarify the physical meaning of the real and imaginary components of energies and currents that arise in the non-Hermitian framework. In non-Hermitian lattice systems, physical observables evaluated within the biorthogonal framework generally acquire complex values, reflecting the intrinsically non-unitary nature of the dynamics. Accordingly, energy eigenvalues take the form $E_{n}=E_{n}^{R}+iE_{n}^{I}$, where the real part $E_{n}^{R}$ determines the spectral position of the state and governs equilibrium-like response properties such as flux sensitivity and state occupation, while the imaginary part $E_{n}^{I}$ encodes the exponential growth or decay rate of the corresponding mode arising from non-reciprocity, asymmetric hopping, or effective gain-loss processes [33, 34, 44]. Persistent charge and spin currents, defined either via the continuity equation or through derivatives of the many-body energy with respect to flux, may likewise possess complex expectation values in non-Hermitian systems. The real part of the persistent charge (spin) current corresponds to the net circulating charge (spin) flow and represents the natural generalization of the Hermitian persistent current, which can in principle be accessed experimentally through magnetic or transport measurements. In contrast, the imaginary part of the current reflects non-conservative probability or spin flow associated with the anti-Hermitian nature of the lattice current operator and has no direct Hermitian analogue. While these imaginary components do not correspond to directly measurable steady-state observables, they carry essential dynamical information and can be indirectly inferred from time-dependent evolution, decay or amplification rates, linewidths, or imbalance of probability flow [45, 46]. Thus, the appearance of imaginary parts in energies and currents is a genuine physical feature of non-Hermitian transport rather than an artifact of the biorthogonal formalism.
@@ II.1 Schematic and model Hamiltonian
II.2 Persistent charge current
The charge current operator is defined as
$\hat{J}=\frac{e\dot{X}}{2Na},\quad\dot{X}=\frac{2\pi}{ih}[X,H],$ (2)
where $e$ is the electronic charge, $a$ is the lattice spacing, and $2N$ is the total number of sites. $X$ and $\dot{X}$ are the position and velocity operators, respectively. In the non-Hermitian case, the current for the $n$th eigenstate is evaluated in the biorthogonal basis (a discussion of biorthogonal basis is given in Appendix A) through the operation as
where $\lvert\psi_{n}^{R}\rangle$ and $\lvert\psi_{n}^{L}\rangle$ denote the right and left eigenvectors, respectively. The explicit derivation of $\hat{J}$ and its evaluation in terms of the expansion coefficients of $\lvert\psi_{n}^{R/L}\rangle$ is provided in Appendix B.
II.3 Persistent spin current
The spin current operator is defined as
The explicit operator forms of $I_{x}$, $I_{y}$, and $I_{z}$, together with their evaluation in terms of the expansion coefficients of $\lvert\psi_{n}^{R/L}\rangle$, are provided in Appendix C.
III Results and Discussion
[…]
III.1 Dispersion relations
We begin the discussion of our results by analyzing the $k$-space spectrum in detail. By writing the Hamiltonian in $k$-space in the basis
@@ III.1 Dispersion relations
where the superscript indicates bonding (-) or antibonding (++) and the subscript indicates spin orientation along $\mathbf{m}$. The full derivation, starting from the Bloch Hamiltonian and proceeding up to the energy spectrum, is provided in Appendix D.
@@
In the upper panel (Figs. 3(a) and (b)), corresponding to $\lvert t_{1}\rvert=0.75$, the system lies in the topological phase, analogous to the non-Hermitian Hatano-Nelson model [36, 39]. The real part of the spectrum (Fig. 3(a)) exhibits a clear bulk energy gap, characteristic of a topologically nontrivial band structure. The imaginary part (Fig. 3(b)), however, is gapless, the bands are smoothly connected across the Brillouin zone without any separation in their imaginary components.
@@
Figure 6: (Color online.) Real and imaginary persistent charge current as a function of the normalized magnetic flux $\Phi/\pi$. The first, second, and third columns correspond to $t_{1}=0.75$, $1$, and $1.25$, respectively. The number of unit cells is $N=20$, the interdimer hopping amplitude is $t_{2}=1$, the spin-dependent splitting is $\mathpzc{h}=0.5$, and the chemical potential is $\mu=0$.
@@
III.2 Real-space analysis: Persistent charge and spin currents
To develop a deeper physical understanding of these results, we next examine a simplified real-space configuration. Specifically, we consider the case where all magnetic moments are aligned along the $z$-quantization axis, i.e., $\theta=0$ and $\phi=0$. In this limit, the Hamiltonian (Eq. 1) decouples into independent up- and down-spin sectors, $\left(H_{\uparrow}+H_{\downarrow}\right)$. We set the SDS parameter to $\mathpzc{h}=0.5$ and fix the number of unit cells in the Hatano-Nelson ring at $N=8$. The resulting complex energy spectra for the up- and down-spin Hamiltonians are shown as a function of $\Phi$ in Fig. 5, for three representative values of the intradimer hopping amplitude. Since the hopping amplitudes follow an SSH-type dimerization, we examine three characteristic regimes: the topological phase $\lvert t_{1}\rvert<t_{2}$, the critical point $\lvert t_{1}\rvert=t_{2}$, and the trivial phase $\lvert t_{1}\rvert>t_{2}$.
@@ III.2 Real-space analysis: Persistent charge and spin currents
Turning to the magnitudes, we observe the following. In the topological phase (Fig. 9(a)), all components are small, within $\pm 0.06\,$eV, reflecting the weak current response in this regime. At the critical point (Fig. 9(b)), the amplitudes of all three components grow significantly, with $I_{z}$ spanning nearly $\pm 0.5\,$eV. In the trivial phase (Fig. 9(c)), the overall trend resembles the earlier cases but with reduced amplitudes (about $\pm 0.15\,$eV) compared to the critical point.
@@
In terms of magnitudes, the topological phase (Fig. 10(a)) again shows weak response, in close analogy with the $\theta$-variation (Fig. 9(a)). Here $I_{z}$ is pinned near 0.030.03\,eV, while $I_{x}$ and $I_{y}$ oscillate within $\pm 0.05\,$eV. At the critical point (Fig. 10(b)), the magnitudes are strongly enhanced. $I_{z}$ remains fixed at about 0.2-0.2\,eV, and $I_{x}$ and $I_{y}$ extend almost up to $\pm 0.4\,$eV. In the trivial phase (Fig. 10(c)), the magnitudes are reduced once again, paralleling the $\theta$-variation (Fig. 9(c)). Here $I_{z}$ stays constant at about 0.060.06\,eV, while $I_{x}$ and $I_{y}$ oscillate within $\pm 0.12\,$eV. Taken together with the $\phi$-dependence and the $\theta$-variation, these results highlight that the spin persistent current is strongly modulated by the orientation of the magnetic moments and respond anisotropically in spin space.
@@
For completeness, we study the finite-size effect on the persistent spin currents by varying the number of unit cells from 10 to 60, while keeping the chemical potential fixed at $\mu=-0.5$. All other parameters are listed in the caption of Fig. 13, where we plot the three components of the persistent spin current as a function of the total number of sites ($2N$, NN being the number of unit cells). We consider three distinct regimes defined by the intradimer hopping amplitude $t_{1}$, with the $x$, $y$, and $z$ components of the persistent spin current shown in red, blue, and green, respectively. While the absolute magnitudes of the spin-current components depend on microscopic parameters such as the chemical potential, hopping strengths, SDS factor, and moment orientations, a systematic decrease of the spin current with increasing system size is observed in all regimes.
Figure 15: (Color online.) Im[$\left[\text{BSBO}_{z}\right]$ as a function of the AAH modulation strength $W$. Panels correspond to (a) $\lvert t_{1}\rvert=0.75$, (b) $\lvert t_{1}\rvert=1$, and (c) $\lvert t_{1}\rvert=1.25$. The system has $N=20$ unit cells, with intercell hopping $t_{2}=1$ and magnetic flux $\Phi=0.25$. The SDS parameter is $\mathpzc{h}=0.5$, with $\theta=\pi/3$ and $\phi=\pi/9$. The chemical potential is set to $\mu=-0.5$.
Importantly, the rate of this decay exhibits a clear and robust phase dependence. In the topological phase (Fig. 13(a)), the decay is the fastest, and the spin current becomes negligibly small beyond $2N\approx 40$. At the critical point defined by $t_{1}=t_{2}$ (Fig. 13(b)), the spin current still decreases with system size. However, the decay is significantly slower, and all components remain finite even for the largest rings considered ($2N=120$). In the trivial phase (Fig. 13(c)), the decay is intermediate between these two limits. This finite-size trend is fully consistent with our previous work [39], where the persistent charge current exhibits a similar behavior with system dimension.
Physically, the universal slowing of the decay at the critical point originates from the vanishing of dimerization at $t_{1}=t_{2}$, which allows spin-dependent hopping processes to propagate more uniformly around the ring. This enhances long-range biorthogonal spin-bond coherence and suppresses finite-size localization effects. Away from the critical point ($t_{1}\neq t_{2}$), dimerization selectively suppresses either intra- or intercell spin transport, leading to a more rapid decay of the spin current with increasing system size. Moreover, we find that the decay is steepest in the topological phase and moderate in the trivial phase. This can be understood from the fact that the effective hopping strength is larger in the trivial phase than in the topological phase, thereby supporting more extended spin transport and a slower decay compared to the topological regime.
@@
To elucidate the microscopic origin of the disorder dependence of the spin current in the non-Hermitian ferromagnetic HN ring, we analyze a bond-resolved biorthogonal quantity that captures spin-dependent coherence between neighboring lattice sites. In non-Hermitian systems, expectation values of physical observables must be evaluated using biorthogonal left and right eigenstates, and the resulting correlators are, in general, complex quantities [43, 44]. For that, we define the biorthogonal spin-bond overlap (BSBO) on a nearest-neighbor bond $i\rightarrow i+1$ as
$\text{BSBO}^{\alpha,(n)}_{i,i+1}=\langle\psi_{n}^{L}|\bm{c}_{i}^{\dagger}\bm{\sigma}_{\alpha}\bm{c}_{i+1}|\psi_{n}^{R}\rangle,\quad\quad\alpha=x,y,z,$ (7)
where $|\psi_{n}^{L}\rangle$ and $\langle\psi_{n}^{R}|$ are the left and right eigenstates of the non-Hermitian Hamiltonian, $\bm{c}_{i}^{\dagger}$ and $\bm{c}_{j}$ are the creation and annihilation operators, respectively as defined in Eq. 1, and $\bm{\sigma}_{\alpha}$ are the Pauli matrices acting in spin space.
For each occupied eigenstate $n$, summing over all bonds yields the spin-bond overlap per eigenstate
$\text{BSBO}^{\alpha,(n)}=\sum_{\langle i,i+1\rangle}\text{BSBO}^{\alpha,(n)}_{i,i+1}$. (8)
Finally, we compute the average BSBO by averaging the state-resolved BSBO over all occupied states with real part of the eigenenergy below the chemical potential $\mu$, which reads as
$\text{BSBO}_{\alpha}=\dfrac{1}{N_{\text{occ}}}\sum_{\text{Re}\left(E_{n}\right)\leq\mu}\text{BSBO}^{\alpha,(n)},$ (9)
where $N_{\text{occ}}$ denotes the number of occupied states.
The BSBO is generally complex. However, only its imaginary part contributes to the spin current. This follows from the anti-Hermitian structure of the lattice current operator, which is obtained either from the continuity equation or equivalently as the derivative of the Hamiltonian with respect to the Peierls phase. As a result, the current selectively probes the phase-dependent component of the biorthogonal bond correlator, while the real part reflects amplitude-like bond coherence that does not contribute to net transport [44].
Since the disorder dependence of the three spin-current components is qualitatively similar apart from their magnitudes, we focus on the imaginary part of the $z$-component of the biorthogonal spin-bond overlap and plot Im[$\left[\text{BSBO}_{z}\right]$ as a function of disorder strength in Fig. 15. We find that Im[$\left[\text{BSBO}_{z}\right]$ closely follows the disorder dependence of the spin current in all three regimes (topological, critical, and trivial), indicating that it provides a microscopic measure of the spin-dependent phase coherence governing transport in the non-Hermitian ferromagnetic HN ring.
The nonmonotonic disorder dependence of the spin current observed across the three regimes is therefore closely connected to the behavior of Im[$\left[\text{BSBO}_{z}\right]$. It is important to note that for the AAH model with isotropic nearest-neighbor hopping in one dimension, a sharp delocalization-to-localization transition occurs at $W=2t$. Consequently, within the disorder window considered in Figs. 14 and 15, the system predominantly resides in the extended regime. Nevertheless, the BSBO profiles indicate that disorder plays a dual role: at moderate strength, it enhances the biorthogonal spin-bond overlap and thereby promotes spin transport, whereas at stronger disorder it suppresses Im[$\left[\text{BSBO}_{z}\right]$ due to the gradual localization of eigenstates. The distinct disorder responses observed in the three regimes thus reflect differences in how biorthogonal phase coherence is reorganized by disorder, rather than signaling sharp transitions between extended and localized states.
@@ IV Summary
$\bullet$ The real and imaginary persistent charge currents are no longer identical in the presence of the ferromagnetic exchange field at the critical point.
@@
Finally, we note that in the present model the persistent spin current is purely real, reflecting the absence of spin splitting in the imaginary spectrum. It would be interesting to explore more general non-Hermitian settings, such as spin-selective gain and loss or non-reciprocal spin-dependent processes, in which spin-dependent non-Hermiticity may generate an imaginary spin current, an issue left for future work.
Appendix A Biorthogonalization of Eigenstates